# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v2

@@ IV-A AO Algorithm
To overcome the challenges posed by the non-concavity of the objective function in (14a) and the intricate coupling between the receive beamforming vectors $\{\mathbf{v}_{k}\}$ and the RA deflection angle vectors $\{\bm{\theta}_{n}\}$, an AO algorithm is proposed to alternately optimize the receive beamforming and RA deflection angles in an iterative manner for the multi-user system.
1) Receive Beamforming Optimization: For given RA deflection angle matrix $\mathbf{\Theta}$, the channel from user $k$ to the BS modeled in (10) becomes fixed. Accordingly, problem (P1) reduces to (by simplifying $\mathbf{h}_{k}(\bm{\Theta})$ to $\mathbf{h}_{k}$)
@@
The SINR in (13) is a generalized Rayleigh quotient with respect to $\mathbf{v}_{k}$, and the receive SINR for each user can be maximized by the MMSE beamforming [37, 38]. Thus, the optimal solution to problem (P4) can be obtained as
@@
$\mathbf{C}_{k}^{-1}=$ $(\mathbf{I}_{N}+\bar{\mathbf{H}}_{k}\mathbf{P}_{k}\bar{\mathbf{H}}_{k}^{H})^{-1}$
$=$ $\mathbf{I}_{N}-\bar{\mathbf{H}}_{k}(\mathbf{P}_{k}^{-1}+\bar{\mathbf{H}}_{k}^{H}\bar{\mathbf{H}}_{k})^{-1}\bar{\mathbf{H}}_{k}^{H}.$ (35)
2) RA Deflection Angle Optimization: For given receive beamforming matrix $\mathbf{V}$, by introducing the slack optimization variable $\eta$ to denote the minimum SINR, problem (P1) can be written as
@@
The above subproblem is still challenging to solve since constraint (36b) is non-convex and the deflection angles, i.e., $\theta_{\mathrm{z},n}$ and $\theta_{\mathrm{a},n}$, are coupled in the pointing vector $\mathbf{f}(\bm{\theta}_{n})$.
Note that the pointing vector $\mathbf{f}(\bm{\theta}_{n})$ is essentially a unit vector on the unit sphere, and the RA deflection angles mainly affect the channel power gains through the projections $\operatorname{cos}(\epsilon_{k,n})$ and $\operatorname{cos}(\tilde{\epsilon}_{k,n})$ as shown in (6) and (7). For ease of subsequent derivation, we introduce an auxiliary variable $\mathbf{f}_{n}\in\mathbb{R}^{3\times 1}$ with $\|\mathbf{f}_{n}\|=1$ to equivalently replace the influence of deflection angle vector $\bm{\theta}_{n}$ on the pointing vector of RA $n$, i.e., $\mathbf{f}_{n}\triangleq\mathbf{f}(\bm{\theta}_{n})$. Thus, based on (8) and (9), the multi-path channel between user $k$ and RA $n$ can be rewritten as
$\tilde{h}_{k,n}(\mathbf{f}_{n})=\alpha_{k,n}\left(\mathbf{f}_{n}^{T}\vec{\mathbf{q}}_{k,n}\right)^{p}+\sum_{q=1}^{Q}{\beta_{k,n,q}\left(\mathbf{f}_{n}^{T}\vec{\mathbf{u}}_{q,n}\right)^{p}},$ (37)
where $\alpha_{k,n}\triangleq\frac{1}{r_{k,n}}\sqrt{\frac{SG_{0}}{4\pi}}e^{-j\frac{2\pi}{\lambda}r_{k,n}}$ and $\beta_{k,n,q}\triangleq\frac{1}{d_{q,n}t_{k,q}}\frac{\sqrt{SG_{0}\sigma_{q}}}{4\pi}e^{-j\frac{2\pi}{\lambda}(d_{q,n}+t_{k,q})+j\chi_{q}}$. Let $\mathbf{F}\triangleq[\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{N}]$ and $\tilde{\mathbf{h}}_{k}(\mathbf{F})\triangleq[\tilde{h}_{k,1}(\mathbf{f}_{1}),\tilde{h}_{k,2}(\mathbf{f}_{2}),\dots,\tilde{h}_{k,N}(\mathbf{f}_{N})]^{T}$, and problem (P7) can be transformed into
@@
s.t. $\frac{\bar{P}_{k}|\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{k}(\mathbf{F})|^{2}}{\sum_{j\neq k}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{j}(\mathbf{F})|^{2}}+1}\geq\eta,\;\forall k,$ (38b)
$\mathbf{f}_{n}^{T}\mathbf{e}_{3}\geq\operatorname{cos}(\theta_{\mathrm{max}}),\;\forall n,$ (38c)
$\|\mathbf{f}_{n}\|=1,\;\forall n,$ (38d)
where constraint (38c) is equivalent to (14b) and constraint (38d) ensures that $\mathbf{f}_{n}$ is a unit vector.
@@
To deal with the fractional structure on the left hand side of constraint (38b), we take the logarithmic operation on both sides of constraint (38b), resulting in an equivalent form for constraint (38b), i.e.,
$\operatorname{ln}\left(\bar{P}_{k}|\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{k}(\mathbf{F})|^{2}\right)\geq\operatorname{ln}(\eta)+\operatorname{ln}\left(\sum_{j\neq k}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{j}(\mathbf{F})|^{2}}+1\right),$ (39)
@@
which is still difficult to handle since $\tilde{h}_{k,n}(\mathbf{f}_{n})$ in (37) is neither convex nor concave due to the complex coefficients $\alpha_{k,n}$ and $\{\beta_{k,n,q}\}$. To tackle this challenge, we adopt the SCA technique to approximate constraint (39) as a convex constraint and obtain a local optimal solution to problem (P6) in an iterative manner. Without loss of generality, we present the procedure of the $(i+1)$-th iteration and denote the solutions of $\mathbf{F}$ and $\eta$ obtained in the $i$-th iteration by $\mathbf{F}^{(i)}$ and $\eta^{(i)}$, respectively. By using the first-order Taylor expansion at $\{\mathbf{f}_{n}^{(i)}\}$, $|\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{k}(\mathbf{F})|^{2}$ and $\operatorname{ln}\left(\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{j}(\mathbf{F})|^{2}}+1\right)$ in (39) can be respectively linearized as $\Lambda_{k}^{(i+1)}(\mathbf{F})$ and $\Gamma_{k}^{(i+1)}(\mathbf{F})$, shown at the top of the next page, where $\tilde{h}_{k,n}^{{}^{\prime}}\triangleq\frac{\partial\tilde{h}_{k,n}(\mathbf{f}_{n}^{(i)})}{\partial\mathbf{f}_{n}^{(i)}}=\tilde{\alpha}_{k,n}\vec{\mathbf{q}}_{k,n}^{T}+\sum_{q=1}^{Q}{\tilde{\beta}_{k,n,q}\vec{\mathbf{u}}_{q,n}^{T}}$ with $\tilde{\alpha}_{k,n}\triangleq\alpha_{k,n}p((\mathbf{f}_{n}^{(i)})^{T}\vec{\mathbf{q}}_{k,n})^{p-1}$ and $\tilde{\beta}_{k,n,q}\triangleq\beta_{k,n,q}p((\mathbf{f}_{n}^{(i)})^{T}\vec{\mathbf{u}}_{q,n})^{p-1}$. Similarly, an upper bound for $\operatorname{ln}(\eta)$ is obtained as $\Xi^{(i+1)}(\eta)\triangleq\operatorname{ln}(\eta^{(i)})+\frac{\eta}{\eta^{(i)}}-1$ by using its first-order Taylor expansions at $\eta^{(i)}$. In this way, constraint (39) can be approximated by
$\Lambda_{k}^{(i+1)}(\mathbf{F})\triangleq|\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{k}(\mathbf{F}^{(i)})|^{2}+\mathrm{Re}\left\{\left(\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{k}(\mathbf{F}^{(i)})\right)^{\ast}\sum_{n=1}^{N}{v_{k,n}^{\ast}\tilde{h}_{k,n}^{{}^{\prime}}(\mathbf{f}_{n}-\mathbf{f}_{n}^{(i)})}\right\}.$ (40)
@@
$\Gamma_{k}^{(i+1)}(\mathbf{F})\triangleq\operatorname{ln}\left(\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{j}(\mathbf{F}^{(i)})|^{2}}+1\right)+\frac{\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}\mathrm{Re}\left\{\left(\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{j}(\mathbf{F}^{(i)})\right)^{\ast}\sum_{n=1}^{N}{v_{k,n}^{\ast}\tilde{h}_{j,n}^{{}^{\prime}}(\mathbf{f}_{n}-\mathbf{f}_{n}^{(i)})}\right\}}}{\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\tilde{\mathbf{h}}_{j}(\mathbf{F}^{(i)})|^{2}}+1}.$ (41)
@@
Problem (P7) is still non-convex due to the unit constraint for $\mathbf{f}_{n}$. For convenience, we first relax the equality constraint (38d) as $\|\mathbf{f}_{n}\|\leq 1$, yielding the following problem.
@@
It can be verified that problem (P8) is a convex optimization problem, which can be solved by the CVX solver [36]. Note that the optimal value obtained by problem (P8) serves as an upper bound for that of problem (P7) due to the relaxation of the equality constraint (38d).
3) Overall Algorithm: We summarize the proposed AO algorithm in Algorithm 1. Since the optimal objective value $\eta$ is non-decreasing over iterations and must be upper-bounded, Algorithm 1 is guaranteed to converge. The complexity order of Algorithm 1 is $\mathcal{O}\left(L(KN^{3}+N^{3.5}\operatorname{ln}(1/\varepsilon))\right)$, where $L$ denotes the required iteration number for algorithm convergence.
After obtaining the optimal solutions of $\mathbf{V}$ and $\mathbf{F}$ through Algorithm 1, the pointing vector needs to be recovered as a unit vector, i.e., $\mathbf{f}_{n}^{\star}=\frac{\mathbf{f}_{n}}{\|\mathbf{f}_{n}\|}$. Furthermore, the original problem aims to obtain the RA deflection angle matrix $\bm{\Theta}$, and thus an additional step is needed to transform the optimized pointing vector into the desired deflection angles. For RA $n$, based on the obtained pointing vector $\mathbf{f}_{n}^{\star}$, the corresponding RA deflection angle vector $\bm{\theta}_{n}^{\star}$ can be obtained as
$\theta_{\mathrm{z},n}^{\star}$ $=\operatorname{arccos}\left((\mathbf{f}_{n}^{\star})^{T}\mathbf{e}_{3}\right),$ (45a)
[…]
As a result, the optimal minimum SINR $\eta^{\star}$ can be calculated by substituting the obtained $\bm{V}^{\star}$ and $\bm{\Theta}^{\star}$ into (13).
@@ IV-B Two-Stage Algorithm
In this subsection, we propose another low-complexity algorithm, namely the two-stage algorithm, to solve problem (P1) without the need for iteration. Specifically, the deflection angles of all RAs are optimized based on the semidefinite relaxation (SDR) technique in the first stage, and the corresponding beamforming vector is obtained by ZF beamformer in the second stage.
@@
As observed in Section IV-A, the difficulty in optimizing the pointing vectors mainly comes from the exponent with power $p$ in the antenna’s directional gain pattern as shown in (5) and the fractional structure of the SINR. According to the law of power conservation, as the directivity factor $p$ increases, the maximum antenna gain $G_{0}$ in the boresight direction becomes larger and the antenna main lobe becomes narrower. Nevertheless, the variation in parameter $p$ does not change the relative magnitude relationship of radiation power in different directions. As such, we consider a typical value of $p$ to eliminate the exponent structure. Specifically, for the cosine-square gain pattern with $p=1$, the multi-path channel between user $k$ and RA $n$ modeled in (37) is expressed as the following linear combination form,
$\hat{h}_{k,n}=\mathbf{f}_{n}^{T}\mathbf{a}_{k,n},$ (46)
where $\mathbf{a}_{k,n}\triangleq\alpha_{k,n}\vec{\mathbf{q}}_{k,n}+\sum_{q=1}^{Q}{\beta_{k,n,q}\vec{\mathbf{u}}_{q,n}}$.
@@
The ZF receive beamforming is adopted to completely remove the inter-user interference, which requires $N\geq K$. Thus, by applying the ZF receive beamforming, the SINR reduces to an SNR without inter-user interference. For user $k$, the ZF receive beamforming, denoted by $\mathbf{v}_{k}^{\mathrm{ZF}}$, should satisfy $(\mathbf{v}_{k}^{\mathrm{ZF}})^{H}\bar{\mathbf{H}}_{k}=\mathbf{0}_{1\times(K-1)}$ with $\bar{\mathbf{H}}_{k}\triangleq[\mathbf{h}_{1},\dots,\mathbf{h}_{k-1},\mathbf{h}_{k+1},\dots,\mathbf{h}_{K}]$. Therefore, the ZF receive beamforming for user $k$ is expressed as
$\mathbf{v}_{k}^{\mathrm{ZF}}=\frac{(\mathbf{I}_{N}-\bar{\mathbf{H}}_{k}(\bar{\mathbf{H}}_{k}^{H}\bar{\mathbf{H}}_{k})^{-1}\bar{\mathbf{H}}_{k}^{H})\mathbf{h}_{k}}{\|(\mathbf{I}_{N}-\bar{\mathbf{H}}_{k}(\bar{\mathbf{H}}_{k}^{H}\bar{\mathbf{H}}_{k})^{-1}\bar{\mathbf{H}}_{k}^{H})\mathbf{h}_{k}\|},\;\forall k,$ (47)
@@
where $\mathbf{I}_{N}-\bar{\mathbf{H}}_{k}(\bar{\mathbf{H}}_{k}^{H}\bar{\mathbf{H}}_{k})^{-1}\bar{\mathbf{H}}_{k}^{H}$ is the projection matrix into the space orthogonal to the columns of $\bar{\mathbf{H}}_{k}$. By substituting (47) into (13), the resultant SINR for user $k$ with ZF beamforming is given by
$\gamma_{\mathrm{ZF},k}=\bar{P}_{k}\|\mathbf{h}_{k}\|^{2}\left(1-\rho_{\mathrm{ZF},k}\right),$ (48)
@@
where $\rho_{\mathrm{ZF},k}=\frac{\mathbf{h}_{k}^{H}\bar{\mathbf{H}}_{k}(\bar{\mathbf{H}}_{k}^{H}\bar{\mathbf{H}}_{k})^{-1}\bar{\mathbf{H}}_{k}^{H}\mathbf{h}_{k}}{\|\mathbf{h}_{k}\|^{2}}$ with $0\leq\rho_{\mathrm{ZF},k}\leq 1$ denoting the SNR loss factor caused by the cancellation of inter-user interference with ZF beamforming. According to (48), the resultant SINR for user $k$ based on ZF beamforming mainly depends on channel power gain $\|\mathbf{h}_{k}\|^{2}$ when the SNR loss factor $\rho_{\mathrm{ZF},k}$ is given by a reasonable value.
@@
1) First stage: Let $\hat{\mathbf{a}}_{k}\triangleq[\mathbf{a}_{k,1}^{T},\mathbf{a}_{k,2}^{T},\dots,\mathbf{a}_{k,N}^{T}]^{T}\in\mathbb{C}^{3N\times 1}$ and $\hat{\mathbf{f}}\triangleq[\mathbf{f}_{1}^{T},\mathbf{f}_{2}^{T},\dots,\mathbf{f}_{N}^{T}]^{T}\in\mathbb{R}^{3N\times 1}$. Then, we formulate the following problem to optimize the pointing vectors.

### v4

@@ IV-A AO Algorithm
To overcome the challenges posed by the non-concavity of the objective function in (14a) and the intricate coupling between the receive beamforming vectors $\{\mathbf{v}_{k}\}$ and the RAs’ pointing vectors $\{\vec{\mathbf{f}}_{n}\}$, an AO algorithm is proposed to alternately optimize the receive beamforming and RAs’ pointing vectors in an iterative manner for the multi-user system.
1) Receive Beamforming Optimization: For a given RA’s pointing matrix $\mathbf{F}$, the channel from user $k$ to the BS modeled in (10) becomes fixed. Accordingly, problem (P1) reduces to (by simplifying $\mathbf{h}_{k}(\mathbf{F})$ to $\mathbf{h}_{k}$)
@@
The SINR in (36a) is a generalized Rayleigh quotient with respect to $\mathbf{v}_{k}$, and thus the receive SINR for each user can be maximized by the MMSE beamforming [43, 44]. Accordingly, the optimal solution to problem (P4) can be obtained as
@@
$\mathbf{C}_{k}^{-1}=$ $(\mathbf{I}_{N}+\tilde{\mathbf{H}}_{k}\mathbf{P}_{k}\tilde{\mathbf{H}}_{k}^{H})^{-1}$
$=$ $\mathbf{I}_{N}-\tilde{\mathbf{H}}_{k}(\mathbf{P}_{k}^{-1}+\tilde{\mathbf{H}}_{k}^{H}\tilde{\mathbf{H}}_{k})^{-1}\tilde{\mathbf{H}}_{k}^{H},$ (38)
where $\tilde{\mathbf{H}}_{k}\triangleq[\mathbf{h}_{1},\dots,\mathbf{h}_{k-1},\mathbf{h}_{k+1},\dots,\mathbf{h}_{K}]$ and $\mathbf{P}_{k}\triangleq\operatorname{diag}\left(\bar{P}_{1},\dots,\bar{P}_{k-1},\bar{P}_{k+1},\dots,\bar{P}_{K}\right)$.
2) RA Pointing Vector Optimization: For a given receive beamforming matrix $\mathbf{V}$, by introducing a slack optimization variable $\eta$ to denote the minimum SINR, problem (P1) can be written as
@@
Note that the above subproblem is still challenging to solve since constraints (14c) and (39b) are non-convex.
Based on (8) and (9), the multipath channel between user $k$ and RA $n$ can be rewritten as
$h_{k}(\vec{\mathbf{f}}_{n})=\alpha_{k,n}\left(\vec{\mathbf{f}}_{n}^{T}\vec{\mathbf{u}}_{k,n}\right)^{p}+\sum_{q=1}^{Q}{\beta_{k,n,q}\left(\vec{\mathbf{f}}_{n}^{T}\vec{\mathbf{c}}_{q,n}\right)^{p}},$ (40)
where
$\alpha_{k,n}$ $\triangleq\frac{1}{r_{k,n}}\sqrt{\frac{AG_{0}}{4\pi}}e^{-j\frac{2\pi}{\lambda}r_{k,n}},$ (41a)
$\beta_{k,n,q}$ $\triangleq\frac{1}{d_{q,n}t_{k,q}}\frac{\sqrt{AG_{0}\sigma_{q}}}{4\pi}e^{-j\frac{2\pi}{\lambda}(d_{q,n}+t_{k,q})+j\chi_{q}}.$ (41b)
Then, we have $\mathbf{h}_{k}(\mathbf{F})=[h_{k}(\vec{\mathbf{f}}_{1}),h_{k}(\vec{\mathbf{f}}_{2}),\dots,h_{k}(\vec{\mathbf{f}}_{N})]^{T}$, and problem (P5) can be transformed into
@@
s.t. $\frac{\bar{P}_{k}|\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\mathbf{F})|^{2}}{\sum_{j\neq k}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\mathbf{F})|^{2}}+1}\geq\eta,\;\forall k,$ (42b)
$\operatorname{cos}(\theta_{\mathrm{max}})\leq\vec{\mathbf{f}}_{n}^{T}\mathbf{e}_{3}\leq 1,\;\forall n,$ (42c)
($,$ (42d)
where constraint (42c) is equivalent to (14b).
@@
To deal with the fractional structure on the left-hand side of constraint (42b), we take the logarithmic operation on both sides of constraint (42b), resulting in an equivalent form for constraint (42b), i.e.,
$\operatorname{ln}\left(\bar{P}_{k}|\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\mathbf{F})|^{2}\right)\geq\operatorname{ln}(\eta)+\operatorname{ln}\left(\sum_{j\neq k}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\mathbf{F})|^{2}}+1\right),$ (43)
@@
which is still difficult to handle since $h_{k}(\vec{\mathbf{f}}_{n})$ in (40) is neither convex nor concave due to the complex coefficients $\{\alpha_{k,n}\}$ and $\{\beta_{k,n,q}\}$. To tackle this challenge, we adopt the SCA technique to approximate constraint (43) as a convex constraint and obtain a local optimal solution to problem (P6) in an iterative manner. Without loss of generality, we present the procedure of the $(i+1)$-th iteration and denote the solutions of $\mathbf{F}$ and $\eta$ obtained in the $i$-th iteration by $\mathbf{F}^{(i)}$ and $\eta^{(i)}$, respectively. By using the first-order Taylor expansion at $\{\vec{\mathbf{f}}_{n}^{(i)}\}$, $|\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\mathbf{F})|^{2}$ and $\operatorname{ln}\left(\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\mathbf{F})|^{2}}+1\right)$ in (43) can be respectively linearized as $\Lambda_{k}^{(i+1)}(\mathbf{F})$ and $\Gamma_{k}^{(i+1)}(\mathbf{F})$, shown at the top of the next page, where $\mathbf{h}_{k,n}^{{}^{\prime}}\triangleq\frac{\partial h_{k}(\vec{\mathbf{f}}_{n}^{(i)})}{\partial\vec{\mathbf{f}}_{n}^{(i)}}=\tilde{\alpha}_{k,n}\vec{\mathbf{u}}_{k,n}+\sum_{q=1}^{Q}{\tilde{\beta}_{k,n,q}\vec{\mathbf{c}}_{q,n}}$ with $\tilde{\alpha}_{k,n}\triangleq\alpha_{k,n}p((\vec{\mathbf{f}}_{n}^{(i)})^{T}\vec{\mathbf{u}}_{k,n})^{p-1}$ and $\tilde{\beta}_{k,n,q}\triangleq\beta_{k,n,q}p((\vec{\mathbf{f}}_{n}^{(i)})^{T}\vec{\mathbf{c}}_{q,n})^{p-1}$. Similarly, an upper bound for $\operatorname{ln}(\eta)$ is obtained as $\Xi^{(i+1)}(\eta)\triangleq\operatorname{ln}(\eta^{(i)})+\frac{\eta}{\eta^{(i)}}-1$ by using its first-order Taylor expansions at $\eta^{(i)}$. In this way, constraint (43) can be approximated by
$\Lambda_{k}^{(i+1)}(\mathbf{F})\triangleq|\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\vec{\mathbf{f}}^{(i)})|^{2}+\mathrm{Re}\left\{\left(\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\vec{\mathbf{f}}^{(i)})\right)^{\ast}\sum_{n=1}^{N}{v_{k,n}^{\ast}\left(\mathbf{h}_{k,n}^{{}^{\prime}}\right)^{T}(\vec{\mathbf{f}}_{n}-\vec{\mathbf{f}}_{n}^{(i)})}\right\}.$ (44)
@@
$\Gamma_{k}^{(i+1)}(\mathbf{F})\triangleq\operatorname{ln}\left(\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\vec{\mathbf{f}}^{(i)})|^{2}}+1\right)+\frac{\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}\mathrm{Re}\left\{\left(\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\vec{\mathbf{f}}^{(i)})\right)^{\ast}\sum_{n=1}^{N}{v_{k,n}^{\ast}\left(\mathbf{h}_{j,n}^{{}^{\prime}}\right)^{T}(\vec{\mathbf{f}}_{n}-\vec{\mathbf{f}}_{n}^{(i)})}\right\}}}{\sum_{j=1,j\neq k}^{K}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\vec{\mathbf{f}}^{(i)})|^{2}}+1}.$ (45)
@@
However, problem (P7) is still non-convex due to the unit-modulus constraint for $\vec{\mathbf{f}}_{n}$ in (14c). For convenience, we first relax the equality constraint (14c) as $\|\vec{\mathbf{f}}_{n}\|\leq 1$, yielding the following problem.
@@
It can be verified that problem (P8) is a convex optimization problem, which can be solved via the CVX solver [45]. Note that the optimal value obtained by problem (P8) serves as an upper bound for that of problem (P7) due to the relaxation of the equality constraint (14c).
@@
3) Overall Algorithm: Based on the results presented in the previous two subproblems, we propose the AO algorithm for problem (P1) by applying the block coordinate descent (BCD) method in Algorithm 1. Specifically, all optimization variables in the original problem (P1) are partitioned into two blocks, i.e., $\{\mathbf{V},\mathbf{F}\}$. Then, the receive beamforming $\mathbf{V}$ and RA’s pointing matrix $\mathbf{F}$ are alternately optimized, by calculating (37) and solving problem (P8), respectively, while keeping the other block of variables fixed. Furthermore, the obtained solution in each iteration is used as the input for the next iteration. The computational complexity of the proposed AO algorithm mainly depends on the matrix inversion for calculating the receive beamforming and the use of CVX solver to optimize the RA pointing matrix, which are performed in an iterative manner. Therefore, the complexity order of Algorithm 1 is $\mathcal{O}\left(L\left(KN^{3}+N^{3.5}\operatorname{ln}(1/\varepsilon)\right)\right)$, where $L$ and $\varepsilon$ denote the required iteration number and accuracy threshold for algorithm convergence, respectively.
Denote the objective value of problem (P1) based on a feasible solution $\{\mathbf{V},\mathbf{F}\}$ as $\eta(\mathbf{V},\mathbf{F})$. In step 4 of Algorithm 1, for a given $\mathbf{F}^{(i)}$, the optimal solution of (P4) is obtained using the MMSE beamformer, which guarantees that $\eta(\mathbf{V}^{(i)},\mathbf{F}^{(i)})\leq\eta(\mathbf{V}^{(i+1)},\mathbf{F}^{(i)})$. Subsequently, in step 5, given $\mathbf{V}^{(i+1)}$ and $\mathbf{F}^{(i)}$, problem (P8) is solved optimally to yield a high-quality solution for problem (P5), thereby ensuring that $\eta(\mathbf{V}^{(i+1)},\mathbf{F}^{(i)})\leq\eta(\mathbf{V}^{(i+1)},\mathbf{F}^{(i+1)})$. As a result, we have $\eta(\mathbf{V}^{(i)},\mathbf{F}^{(i)})\leq\eta(\mathbf{V}^{(i+1)},\mathbf{F}^{(i+1)})$, which indicates that the objective function is non-decreasing over iterations. Since the objective is upper-bounded by a finite value, the AO algorithm is guaranteed to converge to a stationary point.
Note that Algorithm 1 solves the relaxed problem where the equality constraint $\|\vec{\mathbf{f}}_{n}\|=1$ in the original problem (P1) is relaxed to the inequality constraint $\|\vec{\mathbf{f}}_{n}\|\leq 1$. Thus, in the solution obtained by Algorithm 1, if the pointing vector $\vec{\mathbf{f}}_{n}$ is unit-modulus, i.e., the equality in (14c) holds, then the relaxation is tight and the obtained solution is feasible to problem (P1). Otherwise, the pointing vector needs to be reconstructed as a unit vector based on the solution obtained by Algorithm 1, i.e., $\vec{\mathbf{f}}_{n}^{\star}=\frac{\vec{\mathbf{f}}_{n}}{\|\vec{\mathbf{f}}_{n}\|}$.
@@ IV-B Two-Stage Algorithm
In this subsection, we propose another low-complexity algorithm, namely the two-stage algorithm, to solve problem (P1) without the need for iteration. Specifically, the pointing vectors of all RAs are optimized based on the semidefinite relaxation (SDR) technique in the first stage, and the corresponding beamforming vector is obtained by the ZF beamformer in the second stage.
@@
As observed in Section IV-A, the difficulty in optimizing the pointing vectors mainly comes from the power function structure as shown in (40) and the fractional structure of the SINR. According to the law of power conservation, as the directivity factor $p$ increases, the maximum antenna gain $G_{0}$ in the boresight direction becomes larger and the antenna main lobe becomes narrower. Nevertheless, the variation in parameter $p$ does not change the relative magnitude relationship of radiation power in different directions. As such, we consider a typical value of $p$ to eliminate the power function structure. Specifically, for the cosine-square gain pattern with $p=1$, the multipath channel between user $k$ and RA $n$ modeled in (40) is expressed as the following linear combination form,
$\bar{h}_{k}(\vec{\mathbf{f}}_{n})=\vec{\mathbf{f}}_{n}^{T}\mathbf{m}_{k,n},$ (49)
where $\mathbf{m}_{k,n}\triangleq\alpha_{k,n}\vec{\mathbf{u}}_{k,n}+\sum_{q=1}^{Q}{\beta_{k,n,q}\vec{\mathbf{c}}_{q,n}}$.
@@
The ZF receive beamforming is then adopted to completely remove the inter-user interference, which requires $N\geq K$. Therefore, by applying the ZF receive beamforming, the SINR reduces to an SNR without inter-user interference. For user $k$, the ZF receive beamforming, denoted by $\mathbf{v}_{k}^{\mathrm{ZF}}$, should satisfy $(\mathbf{v}_{k}^{\mathrm{ZF}})^{H}\bar{\mathbf{H}}_{k}=\mathbf{0}_{1\times(K-1)}$, where $\bar{\mathbf{H}}_{k}\triangleq[\bar{\mathbf{h}}_{1},\dots,\bar{\mathbf{h}}_{k-1},\bar{\mathbf{h}}_{k+1},\dots,\bar{\mathbf{h}}_{K}]$ with $\bar{\mathbf{h}}_{k}\triangleq[\bar{h}_{k}(\vec{\mathbf{f}}_{1}),\bar{h}_{k}(\vec{\mathbf{f}}_{2}),\dots,\bar{h}_{k}(\vec{\mathbf{f}}_{N})]^{T}$. Therefore, the ZF receive beamforming for user $k$ is expressed as
$\mathbf{v}_{k}^{\mathrm{ZF}}=\frac{(\mathbf{I}_{N}-\bar{\mathbf{H}}_{k}(\bar{\mathbf{H}}_{k}^{H}\bar{\mathbf{H}}_{k})^{-1}\bar{\mathbf{H}}_{k}^{H})\bar{\mathbf{h}}_{k}}{\|(\mathbf{I}_{N}-\bar{\mathbf{H}}_{k}(\bar{\mathbf{H}}_{k}^{H}\bar{\mathbf{H}}_{k})^{-1}\bar{\mathbf{H}}_{k}^{H})\bar{\mathbf{h}}_{k}\|},\;\forall k,$ (50)
@@
where $\mathbf{I}_{N}-\bar{\mathbf{H}}_{k}(\bar{\mathbf{H}}_{k}^{H}\bar{\mathbf{H}}_{k})^{-1}\bar{\mathbf{H}}_{k}^{H}$ is the projection matrix into the space orthogonal to the columns of $\bar{\mathbf{H}}_{k}$. By substituting (50) into (13), the resultant SNR for user $k$ with ZF beamforming is given by
$\gamma_{\mathrm{ZF},k}=\bar{P}_{k}\|\bar{\mathbf{h}}_{k}\|^{2}\left(1-\rho_{\mathrm{ZF},k}\right),$ (51)
where
$\rho_{\mathrm{ZF},k}=\frac{\bar{\mathbf{h}}_{k}^{H}\bar{\mathbf{H}}_{k}(\bar{\mathbf{H}}_{k}^{H}\bar{\mathbf{H}}_{k})^{-1}\bar{\mathbf{H}}_{k}^{H}\bar{\mathbf{h}}_{k}}{\|\bar{\mathbf{h}}_{k}\|^{2}}$ (52)
@@
with $0\leq\rho_{\mathrm{ZF},k}\leq 1$ denoting the SNR loss factor caused by the cancellation of inter-user interference with ZF beamforming. According to (51), the resultant SNR for user $k$ based on ZF beamforming mainly depends on channel power gain $\|\bar{\mathbf{h}}_{k}\|^{2}$ when the SNR loss factor $\rho_{\mathrm{ZF},k}$ is given by a reasonable value.
@@
1) First stage: Let $\mathbf{m}_{k}\triangleq[\mathbf{m}_{k,1}^{T},\mathbf{m}_{k,2}^{T},\dots,\mathbf{m}_{k,N}^{T}]^{T}\in\mathbb{C}^{3N\times 1}$ and $\mathbf{f}\triangleq[\vec{\mathbf{f}}_{1}^{T},\vec{\mathbf{f}}_{2}^{T},\dots,\vec{\mathbf{f}}_{N}^{T}]^{T}\in\mathbb{R}^{3N\times 1}$. Then, we formulate the following problem to optimize the pointing vectors.